OpenStudy (anonymous):

60 N Force is applied to move an object a distance of 4.0 meters. Calculate the amount of work done in joules

OpenStudy (poofypenguin):

Work = force * distance So, W = 60N * 4.0m W= 240 J

OpenStudy (anonymous):

the formula for work done by any object is work =force * distance where force is force applied on that object and distance is distance travelled by it.Acc to eqn the answer is work=F * S w=60 N * 4m w=240J